OpenStudy (anonymous):

how would you find the surface area of a triangular pyramaid with side lenghts of 4in, 4in, and 7in

OpenStudy (anonymous):

the triangular pyramid is composed of 4 triangles and 1 square. so split them up and calculate it individually, then add the surface area together to get the total of the area. For the triangles: both sides are 4 and base is 7. Split this up into two 90 degree triangles. We have the side, and the base, so we calculate the height by using Pythagorean's theorem. \[a ^{2}*b ^{2}=c^{2}\] 3.5^2 + b^2 = 4^2 b = 1.94 The formula for 90 degree triangle is [base*height]/2. so (3.5*1.94)/2 = 3.395in^2 now multiply this by 8 since there are eight 90 degree triangles. 3.395 * 8 = 27.16 we are left with the square to calculate. the surface area for the square is simply 7 * 7 = 49. we add it all up, 27.16 + 49 = 76.16 in ^2

OpenStudy (anonymous):

sorry, i was thinking of a pyramid not a triangular pyramid. You just need to subtract the surface area of two of the 90 degree triangles. and replace the surface area of the square with a triangle. the triangle will have all sides 7 so use pythagorean's theorem again to calculate the surface area.